OpenStudy (anonymous):

Hi! Could someone explain how to solve this? \[\sqrt[3]{4}\times \sqrt[2]{8}\] thank you! <3

OpenStudy (shadowfiend):

Hm. Is that right or is it the opposite? (\(\sqrt[3]{8} \times \sqrt[2]{4}\) ) ?

OpenStudy (anonymous):

sorry, for some reason, the equation writer won't work. and nope, it's not. :D the answer according to my answer key is 4 sqrt[6]{2), but i don't know how to get it.

OpenStudy (anonymous):

If the problem is typed correctly, use index notation and put in powers of 2: cubrt4 = 4^(1/3) = ((2^2)^(1/3) = 2^(2/3) sqrt8 = ((2^3)^(1/2)) = 2^(3/2) Multiply the two by adding indices to get 2^(7/6) which is not the same as your answer, unfortunately

OpenStudy (anonymous):

i'm really sure that the answer is \[4\sqrt[6]{2}\] i even checked it with my calculator myself. but i just don't know how to get it. :(

OpenStudy (anonymous):

\[\sqrt[3]{4}\sqrt{8} \]Square the expression\[\left(\sqrt[3]{4}\sqrt{8}\right)^2=16\ 2^{1/3} \]Take the square root of the result.\[\sqrt{16\ 2^{1/3}}=4\ 2^{1/6}=4 \sqrt[6]{2} \]

OpenStudy (shadowfiend):

Hero in da heezy. Nice one robtobey.

OpenStudy (anonymous):

Aw shucks. Thanks.